(20 points) Mr. Jones has recently purchased 5 head of cattle (meaning, 5 cows) for $20,00o. Using MACRS depreciation, most livestock are considered to have a 5-year recovery period. Mr. Jones expects to be able to sell the cattle for $5,000 after the last depreciation period. Assuming that Mr. Jones should use a 5-year recovery period for each depreciation method, calculate the depreciation schedule (yearly depreciation amounts) using (a) the straight-line method; (b) the . double-declining balance (DDB) method, and (c) the GDS schedule
Answer #1

(a) SL method

Annual depreciation ($) = (Cost - Salvage value) / Life = (20,000 - 5,000) / 5 = 15,000 / 5 = 3,000

SL depreciation schedule is as follows.

SLM
Year Beginning-of-year Book Value ($) Annual Depreciation ($) End-of-Year Book Value ($)
1 20,000 3,000 17,000
2 17,000 3,000 14,000
3 14,000 3,000 11,000
4 11,000 3,000 8,000
5 8,000 3,000 5,000

(b) DDB Method

SLM depreciation rate = 1/Useful life = 1/5 = 0.2

DDB Depreciation rate = 2 x SLM rate = 2 x 0.2 = 0.4

Depreciation schedule is as follows. This method ignores salvage value.

DDB
Year Beginning-of-year Book Value ($) Depreciation Rate Annual Depreciation ($) End-of-Year Book Value ($)
1 20,000 0.4 8,000 12,000
2 12,000 0.4 4,800 7,200
3 7,200 0.4 2,880 4,320
4 4,320 0.4 1,728 2,592
5 2,592 0.4 1,037 1,555

(c) MACRS method

This method ignores salvage value. MACRS depreciation schedule is as follows.

MACRS
Year Depreciation Base ($) Depreciation Rate Annual Depreciation ($) End-of-Year Book Value ($)
1 20,000 20 4,000 16,000
2 20,000 32 6,400 9,600
3 20,000 19.2 3,840 5,760
4 20,000 11.52 2,304 3,456
5 20,000 11.52 2,304 1,152
6 20,000 5.76 1,152 0